5. An open glass tube is immersed in mercury in such a way that a length of 8 cm extends above the mercury
level. The open end of the tube is then closed and sealed and the tube is raised vertically up by additional 46
cm. What will be length of the air column above mercury in the tube now? (Atmospheric pressure = 76 cm
of Hg)
(1) 38 cm (2) 6 cm
(3) 16 cm (4) 22 cm

Grade:12

2 Answers

Taral Jain
7 years ago
Answer is (3) - 16 cm.I`m giving you hint so think about it.(76) (8) = (54 – x) (76 – x)x = 38 cm , Length of air column = 54 – 38 = 16 cm
Taral Jain
7 years ago
Dear student fing the solution below and ask any doubt.Using isothermal conditions8p1=xp Therefore, p=8p1/xNow,8p1/x+(rho)×g×(54-x)=(rho)×g×76Implies: 8/x+(54-x)/76=1Implies: x=16cm.